OpenStudy (cara1382):

I don't know why I am having such a difficult time with this question, but I am & I asked 3 other people & they can't figure it out either. If anybody can give me some directions here, I'd appreciate it! The manager of a coffee ship has one type of coffee that sells for $9/per lb and another type that sells for $15/lb. The manager wishes to mix 40 lbs of the $15 coffee to get a mixture that will sell for $13/lb. How many lbs of the $9 coffee should be used?

OpenStudy (studygurl14):

I think this involves averages. They want a mixture of the $15 and $9 dollar coffee to average out to $13. So, a simple example would be 2 lbs of $15 and 1 l of $9 would result in a mixture worth $13 (\(\Large\frac{2(15)+1(9)}{2}=13\))

OpenStudy (studygurl14):

So...if the manager already has 40 lbs for the $15, the question is how many pounds of the $9 must he add to get an average of $13

OpenStudy (studygurl14):

Do you know how to find that out?

OpenStudy (cara1382):

No! I'm so lost.

OpenStudy (studygurl14):

What, specifically, don't you understand?

OpenStudy (cara1382):

I got help from my brother in law. So simple yet I never would've figured it out. m=mass of $9 per lb 13(m+40)=9m+(40*15) 13m+520=9m+600 4m=8 m=20lbs Thanks so much for your help!